OpenStudy (anonymous):

solve the equation. x^2 - 8x - 20 = 0

OpenStudy (anonymous):

Factoring. Same deal with last one.

OpenStudy (anonymous):

(x - 8)(x + 2) = 0 x = -2, 8

OpenStudy (anonymous):

wrong applebr...........

OpenStudy (anonymous):

yea i saw that, was wrong...

OpenStudy (anonymous):

whoops...sorry (x - 10)(x + 2) = 0 x = -2, 10

OpenStudy (anonymous):

can u guys be my fan

OpenStudy (anonymous):

thxs apple

OpenStudy (anonymous):

solve the equation. x^2 + 4x - 60 = 0